OpenStudy (j3ss):

Using the standard Normal distribution tables, the area under the standard Normal curve corresponding to Z > –1.62 is a. 0.0044. b. 0.0526. c. 0.9474. d. 0.9956.

OpenStudy (j3ss):

Do I have to do 1-Z or is it only when its "Z<-1.62"

OpenStudy (anonymous):

You would have to do 1-Z if you were looking for the area right of -1.62 or Z > -1.62. For your problem, you are simply converting your Z score to a percentage using a Normal Distribution Table which should be located in the back of your stats book. Hoped this help.

OpenStudy (rachida):

the area under the standard Normal curve corresponding to Z > –1.62= P(z>-1.62) P(Z>-1.62)=1-P(Z<1.62) by definition of standard normal disstribution P(Z>-1.62)= 1-0.94738 P(Z>-1.62)=0.05262